OpenStudy (anonymous):

Line AB contains points A (−2, 6) and B (4, 5). The slope of line AB is

OpenStudy (zzr0ck3r):

Slope of line given by connecting points \((x_0,y_0),(x_1,y_1)\) is given by \[\frac{x_0-x_1}{y_0-y_1}\]

OpenStudy (anonymous):

I dont understand

OpenStudy (imstuck):

That slope formula is actually not correct. It's rise over run.

OpenStudy (imstuck):

\[\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }=m\]

OpenStudy (zzr0ck3r):

woops \(\frac{y_0-y_1}{x_0-x_1}\)

OpenStudy (zzr0ck3r):

soory

OpenStudy (anonymous):

haha thanks

OpenStudy (imstuck):

Your y2 is the y coordinate from the second set of parenthesis, which is 5. The y1 is the y coordinate from the first set of parenthesis which is 6. The x2 coordinate from the second set of parenthesis is 4, and the x2 is from the first set of parenthesis which is -2. Your ratio then would look like this:\[\frac{ 5-6 }{ 4-(-2) }\]Can you do that math and find the slope?

OpenStudy (anonymous):

-1/6 right

OpenStudy (imstuck):

Yep! that's right! Good job! See...now that wasn't bad at all, was it!?